# RegD: Hierarchical Embeddings via Distances over Geometric Regions

### v1

@@ Abstract
Hierarchical data are common in many domains like life sciences and e-commerce, and their embeddings often play a critical role. Although hyperbolic embeddings offer a grounded approach to representing hierarchical structures in low-dimensional spaces, their utility is hindered by optimization difficulties in hyperbolic space and dependence on handcrafted structural constraints. We propose RegD, a novel Euclidean framework that addresses these limitations by representing hierarchical data as geometric regions with two new metrics: (1) depth distance, which preserves the representational power of hyperbolic spaces for hierarchical data, and (2) boundary distance, which explicitly encodes set-inclusion relationships between regions in a general way. Our empirical evaluation on diverse real-world datasets shows consistent performance gains over state-of-the-art methods and demonstrates RegD’s potential for broader applications beyond hierarchy alone tasks.
@@ 1 Introduction
Embedding discrete data into low-dimensional vector spaces has become a cornerstone of modern machine learning. In Natural Language Processing (NLP), seminal works such as word2vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) represent words as vectors to capture intricate linguistic relationships. Similarly, knowledge graph embedding methods (Bordes et al., 2013; Sun et al., 2019; Balazevic et al., 2019) encode entities and relations as vectors with their semantics concerned to facilitate reasoning and prediction.
Our work focuses on embedding hierarchical data into low-dimensional spaces. Such data represents partial orders over sets of elements, denoted as $u\prec v$, where $u$ is a parent of $v$. These partial orders naturally manifest as trees or directed acyclic graphs (DAGs). The ability to effectively embed such hierarchical structures enables crucial operations like inferring sub- or superclasses of concepts and classifying nodes within graphs. These capabilities are essential for various tasks in knowledge management and discovery, particularly towards knowledge bases (Shi et al., 2024; Abboud et al., 2020), ontologies (He et al., 2024a; Chen et al., 2024) and taxonomies (Shen & Han, 2022).
@@
In contrast, the hyperbolic approach takes advantage of the unique geometric properties of hyperbolic spaces – specifically their exponential growth in distance and volume – enabling more effective embeddings of tree-structured data in low dimensions. However, existing hyperbolic methods often rely on specialized constructions informed by theoretical analysis or physical models (Ganea et al., 2018; Yu et al., 2024), limiting their applicability beyond strictly hierarchical data. Additionally, training in hyperbolic space is challenging due to precision issues and the complexities of Riemannian optimization.
In this paper, we propose a flexible framework named RegD for modeling hierarchical data by embedding arbitrary regions in Euclidean spaces. Our framework relies on two novel distance metrics, depth distance and boundary distance, which combine the strengths of both region-based approaches in Euclidean spaces and hyperbolic methods.
  • The depth distance (cf. Section 3) enables our model to achieve similar embedding expressivity to hyperbolic spaces by incorporating the “size” of the regions under consideration. Its computation involves only simple operations, thereby reducing computational complexity and eliminating the need for enhanced numerical precision required by hyperbolic methods.
  • The boundary distance (cf. Section 4) improves the representation of set-inclusion relationships among regions in Euclidean spaces. This allows for better identification of shallower and deeper descendants, thereby capturing hierarchical structures more effectively than traditional region-based approaches.
Combining these two distances, we designed a closed formula (cf. Section 5) for determining the partial order of two concepts based on their embedded regions, which is used in both the training and evaluation processes.
Notably, RegD can be applied to arbitrary regions, including common geometric representations such as balls (hyperspheres) and boxes (hyperrectangles). This generality enables broad applicability across diverse geometric embeddings for various tasks, extending beyond hierarchy alone data to ontologies that include hierarchies and more complex relationships in Description Logic (Baader et al., 2017) (cf. Sections 6.2 and 6.3). Our main contributions are summarized as follows:
@@ 2 Preliminaries and Related Works
A $d$-dimensional manifold (Lee, 2013), denoted $\mathcal{M}$, is a hypersurface embedded in an $n$-dimensional Euclidean space, $\mathbb{R}^{n}$, where $n\geq d$, and locally resembles $\mathbb{R}^{d}$. A Riemannian manifold $\mathcal{M}$ is a manifold equipped with a Riemannian metric, enabling the definition of a distance function $d_{\mathcal{M}}(\mathbf{x},\mathbf{y})$ for $\mathbf{x},\mathbf{y}\in\mathcal{M}$. Hyperbolic space, denoted $\mathbb{H}^{n}$, is a Riemannian manifold with a constant negative curvature of $-k$, where $k>0$ (Lee, 2006). It can be represented using various isometric models, such as the Poincaré half-space model, where the points are defined by the half-space: $U^{n}=\{\mathbf{x}\in\mathbb{R}^{n}:\mathbf{x}_{n}>0\},$ and the hyperbolic distance between $\mathbf{x},\mathbf{y}\in U^{n}$ is given by
$d_{\mathbb{H}}(\mathbf{x},\mathbf{y})=\frac{1}{\sqrt{k}}\,\operatorname{arcosh}\left(1+\frac{\|\mathbf{x}-\mathbf{y}\|_{2}^{2}}{2\mathbf{x}_{n}\mathbf{y}_{n}}\right).$
@@
Region-based methods embed the nodes of a directed acyclic graph (DAG) into geometric regions, such as boxes (Boratko et al., 2021), balls (Suzuki et al., 2019), and cones (Vendrov et al., 2016), capturing hierarchical relationships through set-inclusion between these regions. Training is typically conducted using an inclusion loss, which is often defined in terms of the distance or volume of the regions (Vendrov et al., 2016). However, such loss functions may involve non-smooth operations, like maximization, which have been addressed through probabilistic (Vilnis et al., 2018; Dasgupta et al., 2020) or smooth (Boratko et al., 2021) approximations to improve performance.
Due to the crowding effect in the Euclidean space, region-based methods could exhibit suboptimal performance in low-dimensional spaces or when representing large DAGs. To address this challenge, Suzuki et al. (2019) proposed using balls on Riemannian manifolds, enabling more flexible shapes beyond the canonical balls used in Euclidean space. However, leveraging Riemannian manifolds introduces more complex optimization and training processes, which can struggle with tasks beyond hierarchical structures. This makes the approach more restrictive compared to ours, which offers broader applicability.
While Boratko et al. (2021) demonstrates that box-based embeddings in Euclidean spaces can represent any DAG, their focus is on reconstructing DAGs rather than preserving the graph’s transitive properties. Consequently, this approach does not guarantee the faithful capture of hierarchical semantics unless all transitive closure edges are included in the training data.
@@
Hyperbolic space embeddings were first introduced for modeling hierarchical structures by (Nickel & Kiela, 2017). This approach combines hyperbolic distance with the Euclidean norm to represent hierarchies. However, it does not explicitly capture hierarchical relationships and struggles to model the transitivity inherent in these structures. To address these limitations, EntailmentCones (Ganea et al., 2018) and ShadowCones (Yu et al., 2024) were proposed. These methods enhance hierarchy modeling by constructing specific cones in a hyperbolic space: EntailmentCones introduced closed-form hyperbolic cones, determined by their apex coordinates, to ensure transitivity, while ShadowCones is inspired by the physical interplay of light and shadow.
Despite the representational power of hyperbolic spaces, training in non-Euclidean spaces poses unique challenges. Precision-related issues, for example, often arise near the boundaries of the Poincaré half-space model, necessitating high-precision tensor operations that increase both storage requirements and computation time. Moreover, achieving optimal performance in hyperbolic spaces typically requires specialized training techniques, such as burn-in strategies (Yu et al., 2024), custom initialization methods (Ganea et al., 2018), or gradient descent algorithms adapted to Riemannian manifolds (Bécigneul & Ganea, 2019).
@@
3 Depth Distance: Similarly Distance Consdiering Depth
Unlike the Euclidean space, which is constrained by crowding effects that limit its embedding capacity, the hyperbolic space leverage exponential growth in distance and volume to offer superior embedding capabilities. This property makes the hyperbolic space particularly effective for representing tree-structured data in low-dimensional spaces. Notably, two key distinctions arise between region-based embeddings in the Euclidean space and hyperbolic embeddings:
@@ 3 Depth Distance: Similarly Distance Consdiering Depth
    The hyperbolic space better discriminate different hierarchical layers than the Euclidean space.
@@
    The hyperbolic space enable the distinct representation of an arbitrary number of child nodes.
    As demonstrated in Figure 1, the box embeddings in the Euclidean space face inherent limitations when representing multiple children of a node such as animal. As the number of children increases, their corresponding boxes must cluster within $B_{\textit{animal}}$, leading to crowding. In contrast, the hyperbolic space can accommodate an arbitrary number of children while maintaining distinct separations between them. This capability arises from the exponential growth of distance near the boundary of the hyperbolic space, which allows unlimited child nodes to be positioned distinctly by placing them progressively closer to the boundary while preserving meaningful distances between them.
In the following sections, we introduce the depth distance for regions in Euclidean space which considers the “size” of regions. We demonstrate that this measure exhibits advantages similar to those of hyperbolic spaces discussed above (Theorem 1), while maintaining a simpler structure that facilitates implementation. Furthermore, we prove that our depth distance can be equivalent to hyperbolic distance when using specific nonlinear functions, and even with simple polynomial functions, it preserves key properties of hyperbolic geometry (Propositions 1, 2).
3.1 Construction
The depth distance serves as a similarity measure that quantifies the relationship between objects considering their hierarchical depth. As we use set-inclusion relations to model the hierarchy, this hierarchical depth can be represented through the size of the regions, such as their volumes or diameters. Formally, the depth distance is defined as follows, where the size of the regions is represented by a function $f(\textit{reg})$:
Definition 1 (Depth Distance).
Let $\mathcal{R}$ be a collection of regions in the $n$-dimensional Euclidean space $\mathbb{R}^{n}$, where each region $\textit{reg}\in\mathcal{R}$ is characterized by an $m$-dimensional parameter $\textit{par}(\textit{reg})\in\mathbb{R}^{m}$. The depth-aware similarity distance between two regions $\textit{reg}_{1},\textit{reg}_{2}\in\mathcal{R}$, is defined as:
$d_{\text{dep}}(\textit{reg}_{1},\textit{reg}_{2})=g\left(\frac{\|\textit{par}(\textit{reg}_{1})-\textit{par}(\textit{reg}_{2})\|_{p}^{p}}{f(\textit{reg}_{1})f(\textit{reg}_{2})}\right)$ (1)
where $\|\cdot\|_{p}$ is the $p$-norm (i.e., $||\mathbf{x}||_{p}=(\sum_{i}\mathbf{x}_{i}^{p})^{1/p}$), and:
@@ 3.1 Construction
  • $f:\mathcal{R}\to\mathbb{R}_{>0}$ is a function that measures the size of regions. It satisfies: $\lim_{\textit{reg}\to\emptyset}f(\textit{reg})=0.$ [^1] Here, $\textit{reg}\to\emptyset$ indicates that the region reg approaches an empty set, which can be characterized by its volume or other relevant measures.
We require $f$ and $g$ to have non-negative values to ensure the depth distance is non-negative. Additionally, we stipulate that $\lim_{\textit{reg}\to\emptyset}f(\textit{reg})=0$ to guarantee that as a region shrinks to an empty set, the distance between this object and others can approach infinity. This setting emulates the beneficial properties of the hyperbolic space, where the distance between two points can grow rapidly as they approach the boundary of the space (i.e., $\mathbf{x}_{n}=0$ in the Poincaré half-space model). In our context, the boundary of the space $\mathcal{R}$ of a collection of (parametrized) regions in the Euclidean space is the empty set. By selecting an appropriate function $f$, we can control the rate at which the distance between two objects increases as they approach this boundary.
@@
Consider ball defined by a center $\mathbf{c}\in\mathbb{R}^{n}$ and a radius $r>0$ as:
$\textit{ball}(\mathbf{c},r)=\{\mathbf{x}\mid\|\mathbf{x}-\mathbf{c}\|\leq r\}.$
Given two such balls, $\textit{ball}_{1}$ and $\textit{ball}_{2}$, by setting $g(x)=x$ and $f(\textit{ball})=h(\textit{ball})\cdot r$, where $h$ is an arbitrary function defined by the parameters of balls, we obtain their depth distance as:
$d_{\text{dep}}(\textit{ball}_{1},\textit{ball}_{2})=\frac{||\mathbf{c}_{1}-\mathbf{c}_{2}||_{p}^{p}+|r_{1}-r_{2}|^{p}}{(h(\textit{ball}_{1})\cdot r_{1})\cdot(h(\textit{ball}_{2})\cdot r_{2})}.$ (2)
@@
Consider box defined by a center $\mathbf{c}\in\mathbb{R}^{n}$ and an offset $\mathbf{o}\in\mathbb{R}^{n}_{>0}$ as follows:
$\textit{box}(\mathbf{c},\mathbf{o})=\{\mathbf{x}\in\mathbb{R}^{n}\mid\mathbf{c}-\mathbf{o}\leq\mathbf{x}\leq\mathbf{c}+\mathbf{o}\}.$
Given two such boxes, $\textit{box}_{1}$ and $\textit{box}_{2}$, by setting $g(x)=x$ and $f(\textit{box})=h(\textit{box})\cdot\|\mathbf{o}\|$, we obtain the depth distance as:
$d_{\text{dep}}(\textit{box}_{1},\textit{box}_{2})=\frac{||\mathbf{c}_{1}-\mathbf{c}_{2}||_{p}^{p}+||\mathbf{o}_{1}-\mathbf{o}_{2}||_{p}^{p}}{(h(\textit{box}_{1})\cdot\|\mathbf{o}_{1}\|)\cdot(h(\textit{box}_{2})\cdot\|\mathbf{o}_{2}\|)}.$ (3)
The following result highlights that the depth distance exhibits properties analogous to those of the hyperbolic distance discussed earlier in this section. Specifically, the depth distance (1) effectively distinguishes concepts across different layers of the hierarchy with an arbitrary separation gap, denoted by $\Delta$ in item 1 of the following Theorem 1, and (2) distinctly represents an arbitrary number $n$ of children within a single parent by a distance greater than $M$, as demonstrated in item 2 of the theorem.
@@
Consider the region space $\mathcal{B}^{n}$ consisting of balls in $\mathbb{R}^{n}$, with the depth distance defined in Example 1. The following properties hold:
@@
    For any $\textit{ball}_{1},\textit{ball}_{2}\in\mathcal{B}^{n}$ and any $\Delta>0$, there exists a positive constant $r_{0}$ such that for any $\textit{ball}(\mathbf{c}^{\prime},r^{\prime})\subseteq\textit{ball}_{2}$, if $r^{\prime}\leq r_{0}$, then
    $d_{\text{dep}}(\textit{ball}_{1},\textit{ball}(\mathbf{c}^{\prime},r^{\prime}))>d_{\text{dep}}(\textit{ball}_{1},\textit{ball}_{2})+\Delta.$
@@
    For any $\textit{ball}\in\mathcal{B}^{n}$, any integer $n$ and any $M>0$, there exist subsets $_{1},\ldots,\textit{ball}_{n}\subseteq\textit{ball}$ ball such that
    $d_{\text{dep}}(\textit{ball}_{i},\textit{ball}_{j})>M$
    for any distinct $i,j\in\{1,\ldots,n\}$.
@@
Notably, the results established in Theorem 1 naturally generalize to other parameterized regions, with proofs following analogous arguments. Here, we restrict our analysis to boxes and balls due to their well-defined geometric parameters, which allow for clear mathematical formulation and analysis.
3.2 Comparasion with the Hyperbolic Distance
By choosing appropriate functions $f$ and $g$, we can construct a region space that is equivalent to the hyperbolic space, as demonstrated by the following theorem.
@@ 3.2 Comparasion with the Hyperbolic Distance
Let $\mathbb{H}^{n+1}$ be the hyperbolic space with curvature $-1$. Assume the ball space $\mathcal{B}^{n}$ is parameterized as in Example 1, and equipped with the depth distance defined in Equation (1). Then the map
$F\colon\quad\quad\mathcal{B}^{n}$ $\to\mathbb{H}^{n+1}$
$\textit{ball}(\mathbf{c},r)$ $\mapsto[\mathbf{c}:r]$
is an bijective isometry when $p=2$, $g(x)=\operatorname{arcosh}(x+1)$, and $f(\textit{ball}(\mathbf{c},r))=\sqrt{2}\,r$.
Remark 1.
The mapping $F$ above can be generalized to arbitrary regions. For example, we could extend $F$ by replacing the center point and radius of balls with the center of gravity and diameters of any region. However, in this case, the function $F$ might be non-injective, as distinct regions may share the same center of gravity and diameters. Therefore, the resulting region space serves more as an extension rather than an equivalence of hyperbolic space.
Moreover, the following result demonstrates that even with simple linear function $g(\cdot)$, our depth distance retains the ability to capture the hyperbolic structure. Specifically, there exists a map from the region space to the hyperbolic space that preserves the order of hyperbolic distances for all pairs of points, as stated below.
@@
Following Proposition 1, let the depth distance $d_{\text{dep}}$ be redefined by replacing $g$ to $g(x)=k\cdot x+b\,(k>0)$. Then the map $F$ retains the following monotonicity property: for any points $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\mathbf{x}_{4}\in\mathbb{H}^{n}$,
if and only if

### v2

# Achieving Hyperbolic-Like Expressiveness with Arbitrary Euclidean Regions: A New Approach to Hierarchical Embeddings

@@ Abstract
Hierarchical data is common in many domains like life sciences and e-commerce, and its embeddings often play a critical role. While hyperbolic embeddings offer a theoretically grounded approach to representing hierarchies in low-dimensional spaces, current methods often rely on specific geometric constructs as embedding candidates. This reliance limits their generalizability and makes it difficult to integrate with techniques that model semantic relationships beyond pure hierarchies, such as ontology embeddings. In this paper, we present RegD, a flexible Euclidean framework that supports the use of arbitrary geometric regions—such as boxes and balls—as embedding representations. Although RegD operates entirely in Euclidean space, we formally prove that it achieves hyperbolic-like expressiveness by incorporating a depth-based dissimilarity between regions, enabling it to emulate key properties of hyperbolic geometry, including exponential growth. Our empirical evaluation on diverse real-world datasets shows consistent performance gains over state-of-the-art methods and demonstrates RegD’s potential for broader applications such as the ontology embedding task that goes beyond hierarchy.
@@ 1 Introduction
Embedding discrete data into low-dimensional vector spaces has become a cornerstone of modern machine learning. In Natural Language Processing (NLP), seminal works such as word2vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) represent words as vectors to capture intricate linguistic relationships. Similarly, knowledge graph embedding methods Bordes et al. (2013); Sun et al. (2019); Balazevic et al. (2019) encode entities and relations as vectors with their semantics concerned to facilitate reasoning and prediction.
Our work focuses on embedding hierarchical data into low-dimensional spaces. Such data represents partial orders over sets of elements, denoted as $u\prec v$, where $u$ is a parent of $v$. These partial orders naturally manifest as trees or directed acyclic graphs (DAGs). The ability to effectively embed such hierarchical structures enables crucial operations like inferring sub- or superclasses of concepts and classifying nodes within graphs. These capabilities are essential for various tasks in knowledge management and discovery, particularly towards knowledge bases Shi et al. (2024); Abboud et al. (2020), ontologies He et al. (2024a); Chen et al. (2024) and taxonomies Shen & Han (2022).
@@
In contrast, the hyperbolic approach takes advantage of the unique geometric properties of hyperbolic spaces—specifically their exponential growth in distance and volume—enabling more effective embeddings of tree-structured data in low dimensions. However, as shown in Table 1, existing hyperbolic methods often rely on specialized constructed objects as embedding candidates Ganea et al. (2018b); Yu et al. (2024), limiting their generalizability to data that encodes richer semantics beyond hierarchy.
In this paper, we propose a flexible framework named RegD for modeling hierarchical data by embedding arbitrary regions in Euclidean spaces. Our framework relies on two novel dissimilarity metrics, depth dissimilarity $d_{\text{dep}}$ and boundary dissimilarity $d_{\text{bd}}$, which combine the strengths of both region-based approaches in Euclidean spaces and hyperbolic methods. The depth dissimilarity (cf. Section 3.1) enables our model to achieve embedding expressiveness comparable to that of hyperbolic spaces by incorporating the “size” of the regions under consideration (cf. Theorem 1 and Propositions 1 and 3). The boundary dissimilarity (cf. Section 3.2) improves the representation of set-inclusion relationships among regions in Euclidean spaces. This allows for better identification of shallower and deeper descendants, thereby capturing hierarchical structures more effectively than traditional region-based approaches (cf. Proposition 2). Note that these dissimilarities may be negative or fail to satisfy the triangle inequality, and thus are generally not strict metric distances.
Notably, RegD can be applied to arbitrary regions, including common geometric representations such as balls and boxes. This generality enables broad applicability across diverse geometric embeddings for various tasks, extending beyond hierarchy alone data to ontologies that include hierarchies and more complex relationships in Description Logic Baader et al. (2017) (cf. Sections 4.2 and 4.3). Our main contributions are summarized as follows:
@@
For brevity, all proofs of theorems and properties are provided in Appendix A.
Method Embeddings of a node $u$ Energy/score function for $u\prec v$
Poincaré Nickel & Kiela (2017) Points $\mathbf{u}\in\mathbb{H}^{d}$ $(1+\alpha(\|\mathbf{u}\|-\|\mathbf{v}\|))d_{k}(\mathbf{u},\mathbf{v})$
EntailmentCones Cones with apex $\mathbf{u}\in\mathbb{H}^{d}$ and Angle-based function
Ganea et al. (2018b) angle $\theta_{\mathbf{u}}=\arcsin(\frac{(1-\|\mathbf{u}\|^{2})}{\|\mathbf{u}\|})$
ShadowCones Cones with apex $\mathbf{u}\in\mathbb{H}^{d}$ and $d_{\mathbb{H}}(\mathbf{u},\mathbf{v})$ or
Yu et al. (2024) angle $\theta_{\mathbf{u}}=\arctan\sinh(\sqrt{k}r)$ specific boundary distance
RegD (ours) arbitrary regions in $\mathbb{R}^{d}$ $d_{\text{bd}}(\textit{reg}_{u},\textit{reg}_{v})+\lambda\cdot d_{\text{dep}}(\textit{reg}_{u},\textit{reg}_{v})$
(e.g., balls, boxes)
Table 1: Comparison with hyperbolic methods in $\mathbb{H}^{d}$ (curvature k-k, distance $d_{k}$). $\alpha$, $r$ and $\lambda$ are predefined hyperparameters. Only the umbral half-space model of ShadowCones is shown, with boundary distance defined in Eq. 6.
@@ 2 Preliminaries and Related Works
A $d$-dimensional manifold Lee (2013), denoted $\mathcal{M}$, is a hypersurface embedded in an $n$-dimensional Euclidean space, $\mathbb{R}^{n}$, where $n\geq d$, and locally resembles $\mathbb{R}^{d}$. A Riemannian manifold $\mathcal{M}$ is a manifold equipped with a Riemannian metric, enabling the definition of a distance function $d_{\mathcal{M}}(\mathbf{x},\mathbf{y})$ for $\mathbf{x},\mathbf{y}\in\mathcal{M}$. Hyperbolic space, denoted $\mathbb{H}^{n}$, is a Riemannian manifold with a constant negative curvature of k-k, where $k>0$ Lee (2006). It can be represented using various isometric models, such as the Poincaré ball model (see Appendix C) or Poincaré half-space model, where the points are defined by the half-space: $U^{n}=\{\mathbf{x}\in\mathbb{R}^{n}:\mathbf{x}_{n}>0\},$ and the hyperbolic distance between $\mathbf{x},\mathbf{y}\in U^{n}$ is given by $d_{k}(\mathbf{x},\mathbf{y})=\frac{1}{\sqrt{k}}\,\operatorname{arcosh}\left(1+\frac{\|\mathbf{x}-\mathbf{y}\|_{2}^{2}}{2\mathbf{x}_{n}\mathbf{y}_{n}}\right).$
@@
Region-based methods embed the nodes of a directed acyclic graph (DAG) into regions on Euclidean space or Riemannian manifolds, such as boxes Boratko et al. (2021); Zhang et al. (2022), balls Suzuki et al. (2019), and cones Vendrov et al. (2016), capturing hierarchical relationships through set-inclusion between these regions. Training is typically conducted using an inclusion loss, which is often defined in terms of the distance or volume of the regions Vendrov et al. (2016). However, such loss functions may involve non-smooth operations, like maximization, which have been addressed through probabilistic Vilnis et al. (2018); Dasgupta et al. (2020) or smooth Boratko et al. (2021) approximations to improve performance. However, most of these methods suffer from the crowding effect in Euclidean space, which limits their expressiveness compared to our method and hyperbolic-based approaches.
@@
Hyperbolic space embeddings Nickel & Kiela (2017); Sonthalia & Gilbert (2020); Sala et al. (2018) have been introduced to model hierarchical structures by leveraging favorable properties of hyperbolic space, such as its exponential growth. However, these methods primarily rely on pointwise distances to encode hierarchies, which is less intuitive for representing transitivity properties. In contrast, EntailmentCones Ganea et al. (2018b), and ShadowCones propose using regions in hyperbolic space that capture hierarchies through set inclusion, which can be regarded as region-based methods in hyperbolic spaces. Specifically, EntailmentCones introduces closed-form hyperbolic cones, defined by their apex coordinates, while ShadowCones use different cones are inspired by the physical interplay of light and shadow. However, as shown in Table 1, these regions all take specific forms, limiting their generalizability to other tasks or methods.
@@
3 Method
3.1 Depth Dissimilarity: A Similarity Measure Incorporating Depth
Unlike the Euclidean space, which is constrained by crowding effects that limit its embedding capacity, the hyperbolic space leverages exponential growth in distance and volume to offer superior embedding capabilities. This property makes the hyperbolic space particularly effective for representing tree-structured data in low-dimensional spaces. Notably, two key distinctions arise between region-based embeddings in the Euclidean space and hyperbolic embeddings:
@@ 3.1 Depth Dissimilarity: A Similarity Measure Incorporating Depth
    The hyperbolic space better discriminates different hierarchical layers than the Euclidean space.
@@
    The hyperbolic space enables the distinct representation of an arbitrary number of child nodes.
    As demonstrated in Figure 1, the box embeddings in the Euclidean space face inherent limitations when representing multiple children of a node, such as animal. As the number of children increases, their corresponding boxes must cluster within $B_{\textit{animal}}$, leading to crowding. In contrast, the hyperbolic space can accommodate an arbitrary number of children while maintaining distinct separations between them. This capability arises from the exponential growth of distance near the boundary of the hyperbolic space, which allows unlimited child nodes to be positioned distinctly by placing them progressively closer to the boundary while preserving meaningful distances between them.
In the following sections, we introduce the notion of depth dissimilarity for regions in Euclidean space, which explicitly accounts for their “size.” We show that this measure retains advantages analogous to those of hyperbolic spaces (Theorem 1), while offering a simpler structure that facilitates implementation and potential extensions. Moreover, we establish that depth dissimilarity encompasses hyperbolic distance as a special case (Proposition 1), while providing greater flexibility by preserving key properties of hyperbolic geometry through simple polynomial functions (Proposition 2).
Construction
The depth dissimilarity serves as a similarity measure that quantifies the relationship between objects considering their hierarchical depth. As we use set-inclusion relations to model the hierarchy, this hierarchical depth can be represented through the size of the regions, such as their volumes or diameters. Formally, the depth dissimilarity is defined as follows, where the size of the regions is represented by a function $f(\textit{reg})$:
Definition 1 (Depth Dissimilarity).
Let $\mathcal{R}$ be a collection of parameterized regions [^1] See Appendix A.5 for a formal definition. Examples as boxes and balls are shown in Example 1, 2. in the $n$-dimensional Euclidean space $\mathbb{R}^{n}$, i.e., each region $\textit{reg}\in\mathcal{R}$ is characterized by an $m$-dim parameter $\textit{P}(\textit{reg})\in\mathbb{R}^{m}$. The depth dissimilarity between two regions $\textit{reg}_{1},\textit{reg}_{2}\in\mathcal{R}$, is defined as:
$d_{\text{dep}}(\textit{reg}_{1},\textit{reg}_{2})=g\left(\frac{\|\textit{P}(\textit{reg}_{1})-\textit{P}(\textit{reg}_{2})\|_{p}^{p}}{f(\textit{reg}_{1})f(\textit{reg}_{2})}\right)$ (1)
where $\|\cdot\|_{p}$ is the $p$-norm (i.e., $\mathbf{x}||_{p}=(\sum_{i}\mathbf{x}_{i}^{p})^{1/p}$), and:
@@
[…]
We require $f$ and $g$ to have non-negative values to ensure the depth dissimilarity is non-negative. Additionally, we stipulate that $\lim_{\textit{reg}\to\emptyset}f(\textit{reg})=0$ to guarantee that as a region shrinks to an empty set, the dissimilarity between this object and others can approach infinity. This setting emulates the beneficial properties of the hyperbolic space, where the dissimilarity between two points can grow rapidly as they approach the boundary of the space (i.e., $\mathbf{x}_{n}=0$ in the Poincaré half-space model). In our context, the boundary of the space $\mathcal{R}$ of a collection of (parametrized) regions in the Euclidean space is the empty set. By selecting an appropriate function $f$, we can control the rate at which the dissimilarity between two objects increases as they approach this boundary.
For the remainder of this paper, we restrict our theoretical analysis and implementations to boxes and balls as representative region types for clarity and tractability. It is important to note, however, that the theoretical results established herein naturally extend to other parameterized region representations, with analogous proof techniques applicable. For details, please refer to Appendix A.5.
@@
Let $\textit{ball}(\mathbf{c},r)=\{\mathbf{x}\mid\|\mathbf{x}-\mathbf{c}\|\leq r\}$ be a ball defined by a center $\mathbf{c}\in\mathbb{R}^{n}$ and a radius $r>0$. By setting $g(x)=x$ and $f(\textit{ball})=r$, we obtain a depth dissimilarity of the form:
[…]
@@
Let $\textit{box}(\mathbf{c},\mathbf{o})=\{\mathbf{x}\in\mathbb{R}^{n}\mid\mathbf{c}-\mathbf{o}\leq\mathbf{x}\leq\mathbf{c}+\mathbf{o}\}$ be a box defined by a center $\mathbf{c}\in\mathbb{R}^{n}$ and an offset $\mathbf{o}\in\mathbb{R}^{n}_{>0}$. By setting $g(x)=x$ and $f(\textit{box})=\|\mathbf{o}\|$, we obtain a depth dissimilarity:
[…]
The following result highlights that the depth dissimilarity exhibits properties analogous to those of the hyperbolic distance discussed earlier in this section. Specifically, the depth dissimilarity (1) effectively distinguishes concepts across different layers of the hierarchy with an arbitrary separation gap, denoted by $\Delta$ in item 1 of the following Theorem 1, and (2) distinctly represents an arbitrary number $n$ of children within a single parent by a dissimilarity greater than $M$, as demonstrated in item 2 of the theorem.
@@
Consider the region space $\mathcal{B}^{n}$ consisting of balls in $\mathbb{R}^{n}$, with the depth dissimilarity defined in Example 1. The following properties hold:
@@
    For any $\textit{ball}_{1},\textit{ball}_{2}\in\mathcal{B}^{n}$ and any $\Delta>0$, there exists a positive constant $r_{0}$ such that for any $\textit{ball}(\mathbf{c}^{\prime},r^{\prime})\subseteq\textit{ball}_{2}$, if $r^{\prime}\leq r_{0}$, then $d_{\text{dep}}(\textit{ball}_{1},\textit{ball}(\mathbf{c}^{\prime},r^{\prime}))>d_{\text{dep}}(\textit{ball}_{1},\textit{ball}_{2})+\Delta.$
@@
    For any $\textit{ball}\in\mathcal{B}^{n}$, any integer $n$ and any $M>0$, there exist subsets $\textit{ball}_{1},\ldots,\textit{ball}_{n}\subseteq\textit{ball}$ such that for any distinct $i,j\in\{1,\ldots,n\}$, we have $d_{\text{dep}}(\textit{ball}_{i},\textit{ball}_{j})>M.$
@@
Comparison with the Hyperbolic Distance
The following theorem shows that our depth dissimilarity encompasses hyperbolic distance as a special case, obtained with specific choices of functions $f,g$ in Definition 1.
@@
Let $\mathbb{H}^{n+1}$ be the hyperbolic space with curvature 1-1. Assume the ball space $\mathcal{B}^{n}$ is parameterized as in Example 1, and equipped with the depth dissimilarity defined in Equation (1). Then the map $F\colon\mathcal{B}^{n}\to\mathbb{H}^{n+1}$ defined by $F(\textit{ball}(\mathbf{c},r))=[\mathbf{c}:r]$ is an bijective isometry when $p=2$, $g(x)=\operatorname{arcosh}(x+1)$, and $f(\textit{ball}(\mathbf{c},r))=\sqrt{2}\,r$.
Moreover, the following result demonstrates that even with a simple linear function $g(\cdot)$, our depth dissimilarity retains the ability to capture the hyperbolic structure. Specifically, there exists a map from the region space to the hyperbolic space that preserves the order of hyperbolic distances for all pairs of points, as stated below.
@@
Following Proposition 1, let the depth dissimilarity $d_{\text{dep}}$ be redefined by replacing $g$ to $g(x)=k\cdot x+b\,(k>0)$. Then the map $F$ retains the following monotonicity property: for any points $\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\mathbf{x}_{4}\in\mathbb{H}^{n}$, $d_{\mathbb{H}^{n}}(\mathbf{x}_{1},\mathbf{x}_{2})<d_{\mathbb{H}^{n}}(\mathbf{x}_{3},\mathbf{x}_{4})$ if and only if